OpenStudy (anonymous):

solve for x: (2x-5)^2=121

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

\[(2x-5)^2=121\] \[2x-5=11\] or \[2x-5=-11\] so get two answers

OpenStudy (anonymous):

ohhhh... right

OpenStudy (angela210793):

4x^2-20x+25-121=0 4x^2-20x-96=0 D=b^2-4ac and if D>0 then 

OpenStudy (anonymous):

\[2x-5=11\] \[2x=16\] \[x=8\] or \[2x-5=-11\] \[2x=-6\] \[-3\]